xx Testing the Slab thorn

Testing the Slab thorn

Erik Schnetter schnetter@uni-tuebingen.de

Date

Abstract

SlatTest tests the Slab thorn.

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters 7 Interfaces

General Implements: slabtest

Inherits: slab

Grid Variables

7.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11  gfs  10  x01  compact110
11                 10  gfx  x01 description11grid functions
11                 10  gfy  x01 dimensions113
11                 10  gfz  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11  igfs  10  x01  compact110
11                 10  igfx  x01 description11integer grid functions
11                 10  igfy  x01 dimensions113
11                 10  igfz  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11                 10  x01 timelevels111
11                 10  x01variable type11INT
11  a1s  10  x01  compact110
11                 10  a1x  x01 description111D grid arrays
11                 10  a1y  x01 dimensions111
11                 10  a1z  x01 distribution11DEFAULT
11                 10  x01 group type11ARRAY

11                 10  x01  size1110

11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11  a2s  10  x01  compact110
11                 10  a2x  x01 description112D grid arrays
11                 10  a2y  x01 dimensions112
11                 10  a2z  x01 distribution11DEFAULT
11                 10  x01 group type11ARRAY

11                 10  x01  size1110
11                 10  x  x01  size1110

11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11  a3s  10  x01  compact110
11                 10  a3x  x01 description113D grid arrays
11                 10  a3y  x01 dimensions113
11                 10  a3z  x01 distribution11DEFAULT
11                 10  x01 group type11ARRAY

11                 10  x01  size1110
11                 10  x  x01  size1110
11                 10  x01  size1110

11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11 success 10 success x01 compact110
11                 10  x01 description11Indicate a successful test
11                 10  x01 dimensions110
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01variable type11INT

11

Uses header: Slab.h

8 Schedule This section lists all the variables which are assigned storage by thorn CactusNumerical/SlabTest. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:11x
11success 11x

11x  11x

11

Scheduled Functions

CCTKxINITIAL x slabtestxtest x test slab transfer routines x

11x11Language:11c
11 x 11Storage: 11gfs

11 x 11x  11a1s
11 x 11x  11a2s
11 x 11x  11a3s
11 x 11x  11igfs

11 x 11Type: 11function